OpenStudy (anonymous):

find cot x if sin x cotx csc x=√2

OpenStudy (anonymous):

what is csc x in terms of sin and cos?

OpenStudy (anonymous):

the answer is either 4,2,1,√2

OpenStudy (anonymous):

the last option \[\sqrt{2}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Answer my question why don't you just tel her the answer so she can never do it for herself in the future

OpenStudy (anonymous):

I was about to show her, but then again, You're right about that. My mistake.

OpenStudy (anonymous):

print out this webpage http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\sin(x)\times \cot(x)\times cosec(x)=\sqrt{2}\] \[\csc(x)=\frac{1}{\sin(x)}\]

OpenStudy (anonymous):

\[\sin(x) \times \csc(x) \times \cot(x) = \sqrt{2}\]

OpenStudy (anonymous):

\[\frac{\sin(x)}{\sin(x)}=1\]

OpenStudy (anonymous):

\[1\times \cot(x)=\sqrt{2}\]

OpenStudy (anonymous):

Therefore \[\cot(x)=\sqrt{2}\]

OpenStudy (anonymous):

Thanks